the difference of two numbers is 16. the greater number is 5 less than 4 times the smaller number. find the numbers.
Answer by mananth(16946)

x-y =16.......................1
x=5y-4
x-5y=-4........................2
..
multiply equation 1 by -1
-x+y=-16
add this to equation 2
x-x-5y+y=-4+16
-4y=-20
/-4
y=5
...
plug value of y in equation 1
x-(5)=21
x=21
m.ananth@hotmail.ca